OpenStudy (anonymous):

Determine the limit of lim x approaches 5 sqrt(8x+5)-sqrt(45)/x-5

OpenStudy (anonymous):

\[\frac{(\sqrt{8x+5} - \sqrt{45})(\sqrt{8x + 5} + \sqrt{45}) }{ x-5(\sqrt{8x + 5} + \sqrt{45}) } = \frac{ 8x-40 }{ x-5(\sqrt{8x + 5} + \sqrt{45}) }=\frac{ 8 }{(\sqrt{8x + 5} + \sqrt{45}) }\]

OpenStudy (anonymous):

and then take the limitaion

OpenStudy (anonymous):

So then I'd get 8/90, should I reduce?

OpenStudy (anonymous):

i don't know what the answer is but just reduce yourself..